a coffee barista is mixing 2 types of coffee beans elcheapo coffee beans ($2 per pounds) and premium star buckaroo coffee beans ($8 per pound) she has 12 pounds of star buckaroo beans and wishes to mix in a quantity of elcheapoo beans to get a blend that can be sold ofr $6 per pound.
how many pound of elsheapo coffee should be used to create the mixture?

Answer by mananth(16949)

------------------------ price ---------------- quantity
Buckaroo coffee $8 ----------------12 pounds
Elcheapoo coffee $2----------------x pounds
Mixture $6 ----------------12+ x pounds

8*12+2x =6(12+x)
96+2x=72+6x
2x-6x=72-96
-4x=-24
/-4
x=6 pounds of Elcheapoo coffee

m.ananth@hotmail.ca